# InfiJanice: Joint Analysis and In-situ Correction Engine for Quantization-Induced Math Degradation in Large Language Models

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive performance on complex reasoning benchmarks such as GSM8K, MATH, and AIME. However, the substantial computational demands of these tasks pose significant challenges for real-world deployment. Model quantization has emerged as a promising approach to reduce memory footprint and inference latency by representing weights and activations with lower bit-widths. In this work, we conduct a comprehensive study of mainstream quantization methods(e.g., AWQ, GPTQ, SmoothQuant) on the most popular open-sourced models (e.g., Qwen2.5, LLaMA3 series), and reveal that quantization can degrade mathematical reasoning accuracy by up to 69.81%. To better understand this degradation, we develop an automated assignment and judgment pipeline that qualitatively categorizes failures into four error types and quantitatively identifies the most impacted reasoning capabilities. Building on these findings, we employ an automated data-curation pipeline to construct a compact "Silver Bullet" datasets. Training a quantized model on as few as 332 carefully selected examples for just 3–5 minutes on a single GPU is enough to restore its reasoning accuracy to match that of the full-precision baseline.
@@ 1 Introduction
Transformer-based large language models (LLMs), such as LLaMA Grattafiori et al. (2024), GPT Achiam et al. (2023), and QwenYang et al. (2024), have exhibited exceptional proficiency across diverse complex reasoning tasks, including mathematical competitions Maxwell-Jia (2025), code generation Chen et al. (2021), and logical inference Pan et al. (2023). Despite these notable advancements, it is widely recognized that satisfactory performance on intricate reasoning tasks typically demands significantly larger model parameter counts. However, inference using such large-scale models incurs considerable latency and substantial memory consumption, rendering the deployment of full-precision, ultra-large-scale models impractical in most resource-constrained scenarios.
To better balance resource consumption and model performance during deployment, various model compression techniques—including quantization Yang et al. (2019); Rokh et al. (2023), knowledge distillation Hinton et al. (2015); Gou et al. (2021), and pruning Han et al. (2015) have been actively investigated. These approaches streamline models by modifying redundant structures or reducing data precision formats. Quantization, in particular, transforms floating-point data formats into integer arithmetic, dramatically reducing memory footprint and accelerating inference, especially on specialized or resource-limited hardware. Nonetheless, recent studies indicate that operating LLMs at reduced numerical precision (e.g., INT4) significantly degrades performance on tasks involving complex mathematical reasoning Feng et al. (2024); Liu et al. (2025).
In this study, we present a comprehensive and systematic exploration of the impact of Post-Training Quantization Banner et al. (2019) on the mathematical reasoning capabilities of prominent open-source LLMs across multiple benchmarks, including but not limited to MATH Hendrycks et al. (2021), GSM8K Cobbe et al. (2021), and AIME Maxwell-Jia (2025). By format alignment in the reasoning steps, we conduct an extensive analysis to uncover quantization-induced degradation. Our study meticulously characterizes the decline in complex reasoning performance across different model scales and quantization techniques. This analysis provides crucial insights into model deployment within resource-constrained scenarios and outlines effective strategies for mitigating quantization effects to preserve reasoning abilities. Specifically, our primary contributions are as follows:
  • We applied several mainstream quantization methods—AWQ, GPTQ, and SmoothQuant—to some of the most popular open-source LLMs, including the Qwen2.5, LLaMA-3 series. Our results indicate that quantization leads to varying degrees of degradation in reasoning performance. Specifically, we observe quantization consistently reduces reasoning accuracy, and the degradation intensifies as model scale decreases and task complexity increases.
  • We propose a step-aligned assessment protocol for quantitatively and qualitatively dissecting reasoning errors along four dimensions—conceptual, methodological, executional, and logical—encompassing four distinct error types. This protocol facilitates a granular analysis of quantization impacts and the automated error assessment pipeline achieves an accuracy of 97.2% in analysis 9,908 failure cases.
  • We investigate and implement a recovery approach designed to restore the reasoning capabilities of low-bit quantized models. Through targeted data construction and generation techniques, we curate "Silver Bullet" datasets addressing the specific weaknesses introduced by quantization. Remarkably, fine-tuning the quantized model on as few as 332 curated examples for just 3–5 minutes on a single GPU restores its mathematical-reasoning accuracy to that of the full-precision baseline.
@@ 2.1 Quantization Methods
Our methodology encompasses two complementary quantization approaches: (1) Post-training weight-only compression via AWQ Lin et al. (2024) and GPTQ Frantar et al. (2022), achieving 4-bit precision preservation through adaptive rounding strategies; (2) The SmoothQuant Xiao et al. (2023) framework for joint weight-activation quantization, maintaining 8-bit numerical stability via learned scale migration. This dual-strategy design addresses distinct precision requirements: aggressive weight compression for memory efficiency versus moderate activation quantization for computational robustness. Comprehensive implementation protocols, including gradient-aware quantization grid adaptation and layer-wise sensitivity analysis, are detailed in Appendix A.
@@ 2.2 Reasoning Ability Optimization in Large Language Models
LLMs increasingly demonstrate strong general-purpose reasoning skills, spanning commonsense inference to domain-specific problem solving. Early evidence from Minerva Lewkowycz et al. (2022) shows that scaling models and tailoring data can unlock advanced mathematical competence—one instance of the broader trend that rich intermediate computations boost reasoning fidelity. Prompt-engineering techniques such as Chain-of-Thought Wei et al. (2022) and its code-generating variant Program-of-Thought Chowdhery et al. (2023) further improve multi-step reasoning by encouraging models to decompose tasks into interpretable sub-steps.
Orthogonal to prompting, alignment research pursues systematic post-training refinements. Instruction tuning on diversified task mixtures (FLAN) Wei et al. (2021) and lightweight data-curation pipelines (Alpaca) Taori et al. (2023) make models broadly helpful, while Direct Preference Optimization (DPO) Rafailov et al. (2024) offers sample-efficient preference learning without full RLHF. Reliability has been pushed along two complementary axes: self-consistency voting for answer selection Wang et al. (2022) and process-level supervision with stepwise reward models Lightman et al. (2023b), both grounded in verifiable-reasoning theory Creswell and Shanahan (2022).
@@ 3 Methodology
3.1 Investigating Quantization-Induced Reasoning Failures
@@ 3.1.1 Quantization
We conduct a comprehensive investigation into the effects of quantization techniques, encompasses two complementary quantization approaches: (1) Post-training weight-only compression via AWQ Lin et al. (2024) and GPTQ Frantar et al. (2022), achieving 4-bit weight precision preservation through adaptive rounding strategies and keep the data format of activations in 16-bit; (2) The SmoothQuant Xiao et al. (2023) framework for joint weight-activation quantization, maintaining 8-bit numerical stability via learned scale migration. Through the systematic application of these most popular and wild-use quantization techniques, we provide a rigorous and balanced analysis of the resulting quantized models, offering valuable insights into their performance characteristics and trade-offs. Detailed algorithmic descriptions and mathematical derivations for all three methods are provided in Appendix A.
@@ 3.1.2 Format Alignment Training
We employ LoRA Hu et al. (2021) and QLoRA Dettmers et al. (2024) for full-precision model and quantized model respectively as lightweight adaptation techniques for format alignment. These methods efficiently align knowledge of step-by-step solution formats into the model’s latent space without extensive retraining. This fine-tuning enables us to observe how multi-step reasoning is preserved or altered once the model is quantized, offering deeper insights into any capability loss induced by compression.
For alignment, we utilize the PRM800K dataset Lightman et al. (2023a), which provides 800K step-level correctness annotations from 75K solutions to 12K problems. These annotations supply granular, step-by-step reasoning trajectories, equipping models to separate complex problem-solving processes into well-defined stages. To reinforce this structure, we adopt a consistent system prompt across training and evaluation, ensuring that the boundaries of logical steps and final answers are clearly delineated. This consistent, step-by-step alignment is a necessary foundation for our subsequent qualitative and quantitive analyses of quantization-induced degradation in mathematical reasoning. More details are presented on Appendix B
@@ 3.1.3 Detailed Examination of Reasoning Process
Qualitative Analysis To systematically investigate the underlying reasons for degradation in quantized models, we performed a qualitative error analysis inspired by established categorizations from previous literature Brown et al. (2016), Delastri and Lolang (2023) and Kurudirek et al. (2023), which categorize real world student errors in mathematical problem solving. Building on these frameworks, we conduct a qualitative analysis by classifying model-generated errors into seven fine-grained error types, organized under four high-level categories. The definitions of these error types are detailed as follows:
Quantitative Analysis and Error Assessment Pipeline
1. Expert Model Judgement: For each instance in which a quantized model produces an incorrect answer, we employ a dedicated "expert model" to analyze the error. The expert model is tasked with: (a) identifying the first occurrence of an error, (b) specifying the exact step where the error is introduced, (c) assigning an error category based on a nested classification scheme, and (d) providing an explanation along with a confidence score for its determination.
2. Majority Voting: To curb hallucinations and improve evaluation reliability, we apply a three-stage majority-vote protocol to the outputs of five language models—DeepSeek-R1 (primary), GPT-4o, GPT-4, Qwen-Max, and DeepSeek-V3. Instances of disagreement are flagged for further review, ensuring consistency and minimizing spurious judgments.
  1. Rule 1:
    Unanimous agreement ($\geq 4$ votes). If all four auxiliary models concur with the reference judgment from DeepSeek-R1, the answer is accepted.
  2. Rule 2:
    Simple majority ($=3$ votes). If exactly three auxiliary models concur with DeepSeek-R1, the answer is likewise accepted.
  3. Rule 3:
    Escalation ($<3$ votes). Otherwise, the instance is forwarded to two independent human annotators for arbitration.
[…]
@@ 3.2.3 Capability Restoring Training
To reclaim the reasoning capability lost after low-bit quantization, we fine-tune each quantized model using Direct Preference Optimization (DPO) Rafailov et al. (2024). Given a prompt $x$ and a pair of responses $(y^{+},y^{-})$ where $y^{+}$ is the correct answer and $y^{-}$ is the quantized model’s incorrect answer, $y^{+}$ is prefered to $y^{-}$, DPO maximizes the log-likelihood gap between the two while softly constraining the new policy $\pi_{\theta}$ toward the frozen reference policy $\pi_{\mathrm{ref}}$. The objective is
@@
Figure 2: Error assessment results for full-precision and quantized models. For the full-precision model, we aggregate all problems it answered incorrectly; for each quantized model, we count only those problems that the full-precision model solved correctly but the quantized model failed, enabling comparison of quantization-induced changes across error dimensions.
@@ 4.1 Evaluating Quantization Effects
  • AIME (American Invitational Mathematics Examination) is a high-difficulty math competition designed for advanced middle and high school students with 90 problems (from 2022-2025). In 2024, it was adopted as a benchmark for evaluating mathematical reasoning in LLMs. Its difficulty level significantly surpasses that of GSM8K and MATH, approaching the complexity of the International Mathematical Olympiad (IMO).
@@ 4.2 Error Taxonomy and Its Shift Under Quantization
We next analyzed the subset of problems that full-precision models answered correctly but failed under quantized models. Across all three quantization methods, we observed a noticeable increase in the proportion of Method Errors and Execution Errors, suggesting that quantization predominantly impairs the model’s ability to perform procedural operations and arithmetic execution Feng et al. (2024). Supporting this observation, our case study reveals that quantized models exhibit greater difficulty in handling tasks involving basic arithmetic operations and numerical computation.
@@
A careful case study shows that the drop in Reasoning Errors is largely an artefact of our evaluation protocol. Reasoning-type failures typically emerge late in a solution, but the pipeline records only the first erroneous step. Quantization pushes simpler mistakes to earlier stages, so these early slips mask any subsequent logical flaws. In many instances the overall reasoning chain remains valid, yet a premature numerical inaccuracy still leads to an incorrect final answer. Additional illustrative examples are provided in Appendix D.
Figure 3: Capability Restoration from Direct Preference Optimization (DPO) across three evaluation datasets. Light bars represent accuracy before DPO, while dark bars show performance after DPO. Across all quantization settings, DPO consistently restores reasoning ability, effectively narrowing or eliminating the gap with the full-precision (Vanilla-FP16) baseline.
@@ 4.3 Capability Restoration
To prevent data leakage during evaluation, we report results on MATH-500 Lightman et al. (2023b), a 500-problem set that is disjoint from PRM800K yet mirrors the original MATH benchmark in topic coverage and difficulty. Performance on MATH-500 thus reflects genuine reasoning recovery rather than memorization. We also measure accuracy on GSM8K and MMLU Hendrycks et al. (2020) to assess how well the restored model generalises to other reasoning-intensive tasks. The results are visually presented in Figure 3, with additional details provided in Appendix F.
Table 2: Ablation Study on the Effectiveness of Different Training Settings in Error-Aware DPO for Quantized LLMs. Comparison across CE (Conceptual Error), ME (Method Error), EE (Execution Error), Random, and Non-Step settings on GSM8K, MATH500, and MMLU for LLaMA-3.2-3B and Qwen2.5-3B under AWQ/GPTQ quantization.
@@
ALL 74.3 36.8 60.57 73.01 33 59.9 68.84 38.4 64.8 75.21 40.6 63.63 57.42
CE Only 73.19 35.4 60.42 73.31 31.6 59.98 70.28 35.6 65 75.28 34.6 63.61 56.52
ME Only 73.62 32 60.45 72.4 30.4 59.86 69.6 34.8 65.01 76.27 32.6 63.81 55.90
EE Only 73.39 31.2 60.47 72.63 31.2 59.95 69.29 34.8 64.95 76.12 34 63.84 55.98
Random 73.84 30.9 60.51 72.71 31.2 59.97 70.05 32.5 64.92 74.13 34.4 63.71 55.73
Non-Step 70.74 27.8 60.25 69.45 15 59.94 66.79 29.4 65.01 71.04 22.4 64.1 51.83

### v3

# Quantization Meets Reasoning: Exploring and Mitigating Degradation of Low-Bit LLMs in Mathematical Reasoning

@@ Abstract
[…]
@@ 1 Introduction
Transformer-based large language models (LLMs) such as LLaMA (Grattafiori et al., 2024), GPT (Achiam et al., 2023), and Qwen (Yang et al., 2024) have demonstrated strong performance on complex reasoning tasks, including mathematical competitions (Maxwell-Jia, 2025), code generation (Chen et al., 2021), and logical inference (Pan et al., 2023). Yet attaining reliable accuracy on such tasks typically requires large parameter counts. The resulting inference latency and memory footprint make deploying full-precision, ultra-large models impractical in many resource-constrained scenarios. To balance resource use and accuracy, model compression has been extensively studied, including quantization (Yang et al., 2019; Rokh et al., 2023), knowledge distillation (Hinton et al., 2015; Gou et al., 2021), and pruning (Han et al., 2015). Among these, post-training quantization (PTQ) (Banner et al., 2019) lowers precision to reduce memory and improve throughput, especially on edge hardware. However, recent evidence indicates that low-bit operation (e.g., INT4) can substantially degrade mathematical reasoning (Feng et al., 2024; Liu et al., 2025). This raises two practical questions for deployment: Where does degradation emerge in the reasoning process, and How can it be mitigated while remaining in the low-bit regime?
We study these questions through a systematic exploration of PTQ on widely used open-source model families and benchmarks. Concretely, we evaluate AWQ, GPTQ, and SmoothQuant on Qwen2.5 and LLaMA-3 across GSM8K (Cobbe et al., 2021), MATH (Hendrycks et al., 2021), and AIME (Maxwell-Jia, 2025). Using format-aligned chain-of-thought and step-aligned attribution, we characterize quantization-induced failures across model scales and task difficulty. Two patterns are consistent: (i) PTQ predominantly increases method and execution errors (e.g., algorithm choice, rule application, carry/borrow, division/rounding), rather than high-level conceptual mistakes; and (ii) errors tend to emerge early, with the first vulnerable step flipping and cascading to the final answer. This diagnosis turns degradation into a targeted objective: restore token-level margins where collapse happens first.
Guided by this view, as shown on Figure 1, we adopt a lightweight measure–locate–restore loop that operates directly on the quantized model. We first locate the initial erroneous step, then apply small-scale supervised/preference tuning on a compact “Silver Bullet” set designed to target the observed weaknesses. In our experiments, fine-tuning on as few as 332 curated examples for 3–5 minutes on a single GPU is sufficient to recover the mathematical reasoning accuracy of W4A16 models toward their full-precision baselines, while preserving PTQ’s memory and latency benefits.
We frame the study in the regime most relevant to practice: PTQ rather than quantization-aware training, so as to preserve the efficiency budget and expose unmodified low-bit failure modes. To cover current practice, we include AWQ, GPTQ, and SmoothQuant, which together span weight-only and weight–activation designs. Experiments use Qwen and LLaMA models at 0.5–7B—scales commonly deployed under constraints scenario and edge devices. Although our evidence is drawn from these choices, both the step-aligned measurement and the proposed measure–locate–restore loop intervention are architecture- and quantizer-agnostic by construction. Specifically, our primary contributions are as follows:
  • We build a step-aligned measurement suite and hierarchical error taxonomy that expose a robust PTQ-induced shift toward method and execution errors, with earlier first-step flips, consistent across the most popular models, bit-widths, and benchmarks.
  • We develop an automated chain-of-thought error–analysis pipeline (judge ensemble and light human audit) that attains 97.2% labeling accuracy on 9,908 failure cases, enabling fine-grained, reproducible attribution by error type and first faulty step.
[…]
@@ 2.1 Quantization Methods
Our methodology encompasses two complementary quantization approaches: (1) Post-training weight-only compression via AWQ (Lin et al., 2024) and GPTQ (Frantar et al., 2022), achieving 4-bit precision preservation through adaptive rounding strategies; (2) The SmoothQuant (Xiao et al., 2023) framework for joint weight-activation quantization, maintaining 8-bit numerical stability via learned scale migration. This dual-strategy design addresses distinct precision requirements: aggressive weight compression for memory efficiency versus moderate activation quantization for computational robustness. Comprehensive implementation protocols, including gradient-aware quantization grid adaptation and layer-wise sensitivity analysis, are detailed in Appendix A.
@@ 2.2 Reasoning Ability Optimization in Large Language Models
LLMs increasingly demonstrate strong general‐purpose reasoning skills, spanning commonsense inference to domain-specific problem solving. Early evidence from Minerva (Lewkowycz et al., 2022) shows that scaling models and tailoring data can unlock advanced mathematical competence—one instance of the broader trend that rich intermediate computations boost reasoning fidelity. Prompt-engineering techniques such as Chain-of-Thought (Wei et al., 2022) and its code-generating variant Program-of-Thought (Chowdhery et al., 2023) further improve multi-step reasoning by encouraging models to decompose tasks into interpretable sub-steps.
Orthogonal to prompting, alignment research pursues systematic post-training refinements. Instruction tuning on diversified task mixtures (FLAN) (Wei et al., 2021) and lightweight data-curation pipelines (Alpaca) (Taori et al., 2023) make models broadly helpful, while Direct Preference Optimization (DPO) (Rafailov et al., 2024) offers sample-efficient preference learning without full RLHF. Reliability has been pushed along two complementary axes: self-consistency voting for answer selection (Wang et al., 2022) and process-level supervision with stepwise reward models (Lightman et al., 2023b), both grounded in verifiable-reasoning theory (Creswell & Shanahan, 2022).
@@ 3 Methodology
3.1 Quantization-Induced Degradation: Measurement and Attribution
@@ 3.1.1 Quantization
We conduct a comprehensive investigation into the effects of quantization techniques, encompasses two complementary quantization approaches: (1) Post-training weight-only compression via AWQ (Lin et al., 2024) and GPTQ (Frantar et al., 2022), achieving 4-bit weight precision preservation through adaptive rounding strategies and keep the data format of activations in 16-bit; (2) The SmoothQuant (Xiao et al., 2023) framework for joint weight-activation quantization, maintaining 8-bit numerical stability via learned scale migration. Through the systematic application of these most popular and wild-use quantization techniques, we provide a rigorous and balanced analysis of the resulting quantized models, offering valuable insights into their performance characteristics and trade-offs. Detailed algorithmic descriptions and mathematical derivations for all three methods are provided in Appendix A.
@@ 3.1.2 Format Alignment Training
We employ LoRA (Hu et al., 2021) and QLoRA (Dettmers et al., 2024) for full-precision model and quantized model respectively as lightweight adaptation techniques for format alignment. These methods efficiently align knowledge of step-by-step solution formats into the model’s latent space without extensive retraining. This fine-tuning enables us to observe how multi-step reasoning is preserved or altered once the model is quantized, offering deeper insights into any capability loss induced by compression.
For alignment, we utilize the PRM800K dataset (Lightman et al., 2023a), which provides 800K step-level correctness annotations from 75K solutions to 12K problems. These annotations supply granular, step-by-step reasoning trajectories, equipping models to separate complex problem-solving processes into well-defined stages. To reinforce this structure, we adopt a consistent system prompt across training and evaluation, ensuring that the boundaries of logical steps and final answers are clearly delineated. This consistent, step-by-step alignment is a necessary foundation for our subsequent qualitative and quantitive analyses of quantization-induced degradation in mathematical reasoning. More details are presented on Appendix B
@@ 3.1.3 Detailed Examination of Reasoning Process
Qualitative Analysis To systematically investigate the underlying reasons for degradation in quantized models, we performed a qualitative error analysis inspired by established categorizations from previous literature (Brown et al., 2016), (Delastri & Lolang, 2023) and (Kurudirek et al., 2023), which categorize real world student errors in mathematical problem solving. Building on these frameworks, we conduct a qualitative analysis by classifying model-generated errors into seven fine-grained error types, organized under four high-level categories. The definitions of these error types are detailed as follows:
  • Conceptual Errors arise when the model fundamentally misunderstands the underlying principles or context. This includes misgrasping core theories or overlooking domain-specific constraints (e.g., boundary conditions), leading to distorted problem framing and invalid solutions.
  • Method Errors occur when mathematical methods are misapplied or chosen inappropriately. Typical cases include executing standard algorithms incorrectly, skipping key procedures, or misusing formulae in unsuitable contexts.
  • Execution Errors stem from mistakes in arithmetic or symbolic manipulation, such as faulty calculations, erroneous expansions, or mislabeling variables. These slips compromise intermediate computations and ultimately the final answer.
  • Reasoning Errors reflect flaws in logical flow, where inference steps do not follow coherently or essential links are omitted, creating gaps that render the conclusion unsupported.
@@
Quantitative Analysis and Error Assessment Pipeline
1. Expert Model Judgement: For each instance in which a quantized model produces an incorrect answer, we employ a dedicated “expert model” to analyze the error. The expert model is tasked with: (a) identifying the first occurrence of an error, (b) specifying the exact step where the error is introduced, (c) assigning an error category based on a nested classification scheme, and (d) providing an explanation along with a confidence score for its determination.
2. Majority Voting: To curb hallucinations and improve evaluation reliability, we apply a three-stage majority-vote protocol to the outputs of five language models—DeepSeek-R1 (Guo et al., 2025) (primary), GPT-4o, GPT-4, Qwen-Max (Yang et al., 2025), and DeepSeek-V3 (Liu et al., 2024). Instances of disagreement are flagged for further review, ensuring consistency and minimizing spurious judgments. Rule1-Unanimous agreement: If all four auxiliary models concur with the reference judgment from DeepSeek-R1, the answer is accepted. Rule2-Simple majority: If exactly three auxiliary models concur with DeepSeek-R1, the answer is likewise accepted. Rule3-Escalation: Otherwise, the instance is forwarded to two independent human annotators for arbitration.
This pipeline generally aligns with human error analysis and minimizes misclassifications and inconsistencies. However, we also encountered unexpected interesting scenarios. For instance, when the canonical answer is $\frac${11}{2}" but the quantized model outputs "5.5" the expert model sometimes erroneously concludes that there is "No Error" due to subtle discrepancies in reasoning or formatting even if we do not give this type of judgment. Such findings underscore both the robustness and trustiness of our judge framework.
@@ 3.2.3 Capability Restoring Training
To reclaim the reasoning capability lost after low-bit quantization, we fine-tune each quantized model using Direct Preference Optimization (DPO) (Rafailov et al., 2024). Given a prompt $x$ and a pair of responses $(y^{+},y^{-})$ where $y^{+}$ is the correct answer and $y^{-}$ is the quantized model’s incorrect answer, $y^{+}$ is prefered to $y^{-}$, DPO maximizes the log-likelihood gap between the two while softly constraining the new policy $\pi_{\theta}$ toward the frozen reference policy $\pi_{\mathrm{ref}}$. The objective is
@@
Figure 2: Error assessment results for full‐precision and quantized models. For the full‐precision model, we aggregate all problems it answered incorrectly; for each quantized model, we count only those problems that the full‐precision model solved correctly but the quantized model failed, enabling comparison of quantization‐induced changes across error dimensions.
@@ 4.1 Evaluating Quantization Effects
  • AIME (American Invitational Mathematics Examination) is a high-difficulty International Mathematical Olympiad(IMO) competition designed for advanced middle and high school students with 90 problems (we combine problems from 2022-2025 for a widely evaluation).
@@
Figure 3: Relative capability restoration with our method. Radar values are normalized to each model’s Vanilla-FP16 accuracy on the same benchmark (radius 1.0). Solid = After Restoration, dashed = Before Restoration (AWQ, GPTQ).
@@ 4.2 Error Taxonomy and Its Shift Under Quantization
We next analyzed the subset of problems that full-precision models answered correctly but failed under quantized models. Across all three quantization methods, we observed a noticeable increase in the proportion of Method Errors and Execution Errors, suggesting that quantization predominantly impairs the model’s ability to perform procedural operations and arithmetic execution (Feng et al., 2024). Supporting this observation, our case study reveals that quantized models exhibit greater difficulty in handling tasks involving basic arithmetic operations and numerical computation.
@@
A careful case study shows that the drop in Reasoning Errors is largely an artefact of our evaluation protocol. Reasoning-type failures typically emerge late in a solution, but the pipeline records only the first erroneous step. Quantization pushes simpler mistakes to earlier stages, so these early slips mask any subsequent logical flaws. In many instances the overall reasoning chain remains valid, yet a premature numerical inaccuracy still leads to an incorrect final answer. Absolute accuracies and additional illustrative examples are reported in Appendix D.
@@ 4.3 Capability Restoration
To prevent data leakage during evaluation, we report results on MATH-500 (Lightman et al., 2023b), a 500-problem set that is disjoint from PRM800K yet mirrors the original MATH benchmark in topic coverage and difficulty. Performance on MATH-500 thus reflects genuine reasoning recovery rather than memorization. We also measure accuracy on GSM8K and MMLU (Hendrycks et al., 2020) to assess how well the restored model generalises to other reasoning-intensive tasks. The results are visually presented in Figure 3, with additional details provided in Appendix F.
Table 2: Ablation results on GSM8K, MATH500, and MMLU under AWQ/GPTQ (W4A16). Row labels denote training subsets: ALL-S = all error cases with step-aligned supervision from the first error; CE/ME/$E$-S = only conceptual/method/execution errors with step alignment; Rand-NS = size-matched random sampling without step alignment; ALL-NS = all error cases without step alignment. Numbers are accuracy (%).
@@
ALL-S 74.3 36.8 60.57 73.01 33 59.9 68.84 38.4 64.8 75.21 40.6 63.63 57.42
CE-S 73.19 35.4 60.42 73.31 31.6 59.98 70.28 35.6 65 75.28 34.6 63.61 56.52
ME-S 73.62 32 60.45 72.4 30.4 59.86 69.6 34.8 65.01 76.27 32.6 63.81 55.90
$E$-S 73.39 31.2 60.47 72.63 31.2 59.95 69.29 34.8 64.95 76.12 34 63.84 55.98
Rand-NS 73.84 30.9 60.51 72.71 31.2 59.97 70.05 32.5 64.92 74.13 34.4 63.71 55.73
ALL-NS 70.74 27.8 60.25 69.45 15 59.94 66.79 29.4 65.01 71.04 22.4 64.1 51.83